OpenStudy (anonymous):

HELP 7m − 5 − 13m = 25 show all work please

OpenStudy (anonymous):

7m − 5 − 13m = 25 -6m-5=25 +5 +5 -6m=30 -6m/-6=30/-6 m=-5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome